OpenStudy (anonymous):

calculate the specific heat j/gc for 50.2g sample of a metal that absorbs 355j when tempture increases 35.0c to 57.9c

OpenStudy (julian.pitalua):

Start with the equation for heat absorbed of expelled.... Q= Specific heat x Mass x Temperature change then... Specific heat= Q/mass x temperature change Replacing... \[Specific heat= 355J \div (50.2 g \times (57.9 º C - 35.0 º C)= xxx J / g \times ºC\]